OpenStudy (anonymous):

what is the slope of the line normal to the graph y= x^2/ (3 in the radical thing, not sure what its called) √(3x^2+1) at x =1

zepdrix (zepdrix):

\[\Large\bf\sf y=\frac{x^2}{\sqrt[3]{3x^2+1}},\qquad\qquad\text{at } x=1\]

zepdrix (zepdrix):

That look right?

OpenStudy (anonymous):

yes :)

zepdrix (zepdrix):

We only need the `slope` of the normal line? We don't need to setup an equation for it? Oh that's nice :) So we'll start by finding the slope of the line `tangent` to the curve at x=1. How do we find tangent lines? Remember? :X

OpenStudy (anonymous):

we derive to get slope? sorry;) and then y =mx +b

zepdrix (zepdrix):

Since they didn't ask for the `equation` of the line normal to the curve, we don't have to worry about the y=mx+b part! We'll simply find the m, (which as you said, we'll derive to find it) and then we'll do some magic to find the slope of the normal line.

OpenStudy (anonymous):

lol

zepdrix (zepdrix):

Have you tried taking a derivative yet? This one is a bit of a doozy. Make sure you don't try to simplify.

OpenStudy (anonymous):

no so we use quotient rule

zepdrix (zepdrix):

sounds good.

OpenStudy (anonymous):

ok ill do

zepdrix (zepdrix):

\[\Large\bf\sf y'=\frac{\color{orangered}{(x^2)'}\sqrt[3]{3x^2+1}-x^2\color{orangered}{(\sqrt[3]{3x^2+1})'}}{\left(\sqrt[3]{3x^2+1}\right)^2}\] There's our quotient rule setup, Gotta differentiate the orange parts. Get to it girl! :O

OpenStudy (anonymous):

i finished lol

zepdrix (zepdrix):

Hmm kinda hard to write it all out on here isn't it? :\

OpenStudy (anonymous):

jk still working on it wait im simplifying

zepdrix (zepdrix):

\[\Large\bf\sf y'=\frac{\color{royalblue}{(2x)}\sqrt[3]{3x^2+1}-x^2\cdot\color{royalblue}{6x(3x^2+1)^{-2/3}}}{\left(\sqrt[3]{3x^2+1}\right)^2}\]

zepdrix (zepdrix):

Nooo don't simplify :O that's a waste of time.

zepdrix (zepdrix):

Did you get something like that?

OpenStudy (anonymous):

lol yea i got that

zepdrix (zepdrix):

Woops I forgot my 1/3 in front of the second term. my bad.

zepdrix (zepdrix):

From the power rule on the cube root term.

zepdrix (zepdrix):

\[\Large\bf\sf y'=\frac{\color{royalblue}{(2x)}\sqrt[3]{3x^2+1}-x^2\cdot\color{royalblue}{2x(3x^2+1)^{-2/3}}}{\left(\sqrt[3]{3x^2+1}\right)^2}\]

zepdrix (zepdrix):

Now we'll evaluate it at x=1. This is why we don't want to simplify ahead of time :U Plugging in 1 should shrink it down pretty quickly.

OpenStudy (anonymous):

hold pls i messed up whit the exponent s ill tell u

OpenStudy (anonymous):

ok im their

OpenStudy (anonymous):

so now we just plug in 1 to F(x)

OpenStudy (anonymous):

y'x

zepdrix (zepdrix):

ya find y'(1).

OpenStudy (anonymous):

31/4??

OpenStudy (anonymous):

wait forgot the bottom

OpenStudy (anonymous):

sorry im confusing myself

zepdrix (zepdrix):

Hmm this one is a doozy :\

OpenStudy (anonymous):

yea help me pls

zepdrix (zepdrix):

Mmm let's assume we did our derivative correctly and just see what happens. \[\Large\bf\sf y'=\frac{(2)4^{1/3}-2(4)^{-2/3}}{(4)^{2/3}}\]

zepdrix (zepdrix):

So that's what I got when I plugged 1 in. Were you able to get to that step or so? :o

OpenStudy (anonymous):

i forgot how to convert the square roots to fractions but i got it now

OpenStudy (anonymous):

thank you

zepdrix (zepdrix):

I hope we did that correctly D: Imma try the derivative again.. looks weird..

zepdrix (zepdrix):

Ya if you use your calculator to simplify it down further, you'll end up with,\[\Large\bf\sf y'(1)\approx 0.9449408\]Which is good. That's correct, so our derivative looks good. That's the slope of the line `tangent` to our curve at x=1. Remember how a normal line relates to a tangent line?

zepdrix (zepdrix):

If your teacher wants the answer in exact form, then maybe we should avoid the decimal :d it's up to you.

OpenStudy (anonymous):

i honestly forgot but something simple

zepdrix (zepdrix):

ya simple :d you flip and negate. The slope of the normal line will be the `negative reciprocal` of the slope of the tangent line.

zepdrix (zepdrix):

\[\Large\bf\sf m_{\tan}\quad=\quad -\frac{1}{m_{norm}}\]

OpenStudy (anonymous):

so is .945 the slope of the normal line b/c its one of my answers?

zepdrix (zepdrix):

It's the slope of the tangent line. Or it should be at least :x

zepdrix (zepdrix):

For the slope of the normal line, we need to flip it, and negate it.\[\Large\bf\sf m_{\tan}\approx 0.9449408\] The answer we're looking for, the slope of the normal line,\[\Large\bf\sf m_{norm}\approx -\frac{1}{0.9449408}\]

zepdrix (zepdrix):

Hopefully this will give us one of your answer choices also ^^ lol

zepdrix (zepdrix):

Do you have a fancy calculator you can do that step with? :D

OpenStudy (anonymous):

yea i fgot the ti inspire thank you

OpenStudy (anonymous):

Thank You Zepdrik Goodnight <3

zepdrix (zepdrix):

get the right answer? :O go to bed!